Water flows at a rate of 800 cm³/min in a pipe of varying cross-sectional areas. Calculate the velocity of the water at the section where;

a) the radius is 3.0 cm; and

b) the radius is 5.0 cm


1
Expert's answer
2021-03-14T19:15:38-0400

a)


"C=vA\\800=\pi(3)^2v\\v=28\frac{cm}{min}"

b)


"C=vA\\800=\pi(5)^2v\\v=10\frac{cm}{min}"